i need help solving linear equations i dont remember how to exactly. help!!!
2x-2(2-x)=40
2(2x+4)=-3(2x+1)
12-(3x-2)=2(x-1)+3x
-3(x-2)=5(2x-12)

Answer by oscargut(2103)

a) 2x-2(2-x)=40
2x-4+2x=40
4x-4=40
4x=44
x=11
b) 2(2x+4)=-3(2x+1)
4x+8=-6x-3
10x=-11
x=-1.1

c) 12-(3x-2)=2(x-1)+3x
12-3x+2=2x-2+3x
-8x=-16
x=2

d) -3(x-2)=5(2x-12)
-3x+6=10x-60
-13x=-66
x=66/13